OpenStudy (anonymous):

3x^2 + 4x = 6 solve using the quadratic formula

hartnn (hartnn):

@trippytripp94 Hi, \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\) Compare your quadratic equation with \(ax^2+bx+c=0\) find a,b,c then the two roots of x are: \(\huge{x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

hartnn (hartnn):

before that write 3x2 + 4x = 6 as 3x^2+4x-6=0

OpenStudy (anonymous):

whats the answer?

hartnn (hartnn):

are you here just for answers ?

hartnn (hartnn):

or to learn how to solve ?

OpenStudy (anonymous):

umm..yea

hartnn (hartnn):

yes for ?

OpenStudy (anonymous):

i just want the answer bro

hartnn (hartnn):

read http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

dude this is the internet, nobody cares. Soo you wont give the answer?

hartnn (hartnn):

i don't think anybody here will give u answer. if u want just answer use http://www.wolframalpha.com/

hartnn (hartnn):

just type 3x^2 + 4x = 6 there.